ID |
Date |
Author |
Subject |
Text |
 |
154
|
Tue Nov 22 09:26:04 2016 |
A. Tarifeno-Saldivia | Noise check | Run for checking the noise just after turn
on the crates, without high voltages on the
PREAMPS.
|
|
153
|
Tue Nov 22 08:28:36 2016 |
A. Tarifeno-Saldivia et al | New directory and Config File | A new directory in the IFICdaq computer has
been created for the parasitic measurements
during the Fallon Beam Time.
|
|
152
|
Sun Nov 20 13:02:06 2016 |
Jorge, Alavaro, Jose | DACQ | Gasific-7.0 Short Manual (Preliminary) |
|
151
|
Sun Nov 20 12:28:27 2016 |
Jorge, Alvaro, Jose | DACQ | Updated run protocol |
|
150
|
Fri Nov 11 08:07:18 2016 |
BCR SG NTB | BRIKEN Hit Visualization | These are the files used to create and
visualize the hit patterns of the BRIKEN
detector. |
  |
Draft
|
Thu Nov 10 05:05:54 2016 |
KM | How to convert ridf-file to root-file | 1. Login to briken@ribfana03.riken.jp
2. Check if the ridf-file you want
to convert is in the directory; /home/briken16/rawdata/bigrips/ |
|
148
|
Wed Nov 9 02:02:22 2016 |
Iris | Boron rubber from India | The boron rubber from India is BR-2871, price
US-$ 2700 for 2pc 1m x 2m x 3.3mm.
52% B content, density 1.61 g/cm3 |
  |
Draft
|
Wed Nov 9 01:04:35 2016 |
Iris | Drawing of present setup | This is the setup which was used for the Ni
runs (from Nov. 8th evening on).
|
|
146
|
Wed Nov 9 01:02:53 2016 |
Iris | Elog | From Nov. 9th on we use the Takechi ELOG for
the run information:
http://ribf-exp.riken.jp/elog/RIBF123/ |
|
145
|
Tue Nov 8 17:00:27 2016 |
NTB, CG, AT | Si Delta E | At least during the timeframe in between
~ 21:30 (11/08) - 23:30 (11/09) (161108_2323_717273Ni.root)
we see that the silicon dE have a very low |
|
144
|
Tue Nov 8 15:10:24 2016 |
Many | 717273Ni Run | F7 OUT F11 C-6mm
Start: 00:09
|
|
143
|
Tue Nov 8 14:28:54 2016 |
Many | 717273Ni run | F7 CH2 F11 C-6mm
Total degrader thickness 6mm
|
|
142
|
Tue Nov 8 13:49:58 2016 |
Many | 71,72,73Ni | F7 CH2 F11 Empty
We remove the 8mm Al plate from inside BRIKEN |
|
141
|
Tue Nov 8 12:51:08 2016 |
Many | 717273Ni | After removing the lead wall.
Start: 21:45
|
|
140
|
Tue Nov 8 12:30:33 2016 |
Tain, Ale (who are not on shift) | 71,72,73Ni run | 71,72,73Ni run
We started at approximately 21:00, only online, |
|
139
|
Tue Nov 8 11:51:28 2016 |
Ale | spikes on some spectra | Some spectra show spikes or do not look normal.
This should be checked.
|
|
138
|
Tue Nov 8 08:49:51 2016 |
Tain et al | Move ELOG | We will start to use the internal RIKEN ELOG
It is not clear that we can keep the present
ELOG also updated |
|
137
|
Tue Nov 8 08:31:52 2016 |
Tain et al | 424344Ca run | Sequence of settings:
1) No target at F7 and Empty target at F11
(up to t=1E6ms)
|
|
136
|
Tue Nov 8 06:39:25 2016 |
Tain. Iris | Drawing of present setup |
Attached is the present setup including all |
|
135
|
Tue Nov 8 06:13:11 2016 |
Tain et al | New plastic detector in F11 | There is a new plastic scintillator
after the MUSIC chambers
|
 |